Write the equation of the corresponding states and explain the meaning of the terms involved in it.


1
Expert's answer
2021-04-13T02:07:49-0400

Ideal gas equation of state is PV =RT for one mole of gas

Van der Waals equation of state proposed the following equation to explain the P-V-T behaviour of real gases

The pressure term in ideal gas equation replaced by P + (a/V2)

and Volume term is replaced by (V - b )

Where b is the volume per mole of the gas

(P + a/V2)(V-b) = RT

V is the molar volume of the gas

VC = 3b,. TC = 8a/27Rb.

PC = (RTC/2b)-(a/9b2)

on solving equations of TC and VC we get Van der Waals constants as

a= 27R2TC2/64PC

b= RTC/8PC

TC and PC are critical temperature and critical pressure

a and b are the van der Waals constant

P is the actual pressure of the gas


The principal states that all gases when compared at the same reduced temperature and the reduced pressure, have approximately the same compressibility factor and all deviate from the ideal behaviour to the same extent

Reduced pressure Pr = P/PC

Reduced temperature Tr = T/TC and

Reduced Volume Vr = V/VC

Where V, T and P are the actual Volume, temperature and pressure

(Pr + 3/Vr2)(3Vr - 1) = 8Tr ________(a)

above equation is known as Van der Waals equation of state in reduced formed